OpenStudy (anonymous):

evalute using convolution theorem 1/(s+1)(s+9)^2

OpenStudy (anonymous):

To be honest I've never even come across the convolution theorem. Sorry I can't be of more help! I'll tag a few people of higher level who may be able to assist. @agent0smith @primeralph @dan815

OpenStudy (anonymous):

thanks for ur coperation miss sarahusher>>>...........

OpenStudy (loser66):

@TuringTest

OpenStudy (turingtest):

I suppose the theorem we are supposed to use here is\[\mathcal L^{-1}\{F(s)G(s)\}=(f*g)(t)\]

OpenStudy (turingtest):

so let \[F(s)=\frac1{s+1}\implies f(t)=e^{-t}\]by form 2 on this page http://tutorial.math.lamar.edu/Classes/DE/Laplace_Table.aspx and\[G(s)=\frac1{(s+9)^2}\implies g(t)=e^{-9t}t\]by form 23 so now we need to use the theorem

OpenStudy (turingtest):

\[\large(f*g)(t)=\int_0^tf(t-\tau)g(\tau)d\tau=\int_0^te^{-(t-\tau)}\cdot e^{-9\tau}\tau d\tau\\\large\mathcal L^{-1}\left\{{1\over(s+1)(s+9)^2}\right\}=\int_0^te^{-t-8\tau}\tau d\tau\]which looks doable by integration by parts

OpenStudy (anonymous):

Turing 1stly thanks alot for solving my question..............r u a person or a comunity????

OpenStudy (turingtest):

I'm just a student living in Mexico :)

OpenStudy (anonymous):

WHATS UR NAME???????